OpenStudy (anonymous):

Let (X,d) be a metric space and \[x\in X\], show that \[\left\{ y\in X|d(y,x)>r \right\}\]is open for any\[r \in \mathbb{R} \]

OpenStudy (anonymous):

If question can't load: Let (X,d) be a metric space and \[x\in X\], show that \[\left\{ y\in X|d(y,x)>r \right\}\]is open for any\[r \in \mathbb{R} \]

OpenStudy (purple_pink):

huh?

Parth (parthkohli):

@ikram002p I think you're the only qualified one for this.

Parth (parthkohli):

Oh, hmm. This question is pretty intuitive actually. Here is what Wikipedia says: \[B(x;r) = \{y\in M:d(x,y) < r\}\]Explicitly, a subset U of M is called open if for every x in U there exists an r > 0 such that B(x;r) is contained in U. This is all I know about metric spaces but I'm going to try to help you since you paid for this and Ikram isn't here yet.

Parth (parthkohli):

Oh, so you're asking if the set you posted is an "open" set in that way? And ikram is here so I guess I'll leave.

OpenStudy (anonymous):

Could you send me the link for that site?

OpenStudy (ikram002p):

ok i'm here now :D. let me conjuring my mind up...

Parth (parthkohli):

I guess it makes sense because the set is a part of the standard topology on \(\mathbb R^2\) haha. I'll stop.

OpenStudy (ikram002p):

@Tommynaiter you need to show these stuff (familiar with them?) \(\Large \tau=\left\{ y\in X|d(y,x)>r \right\}\) the trivial step i- \(\varnothing,X \in \tau \) ii- \(\Large U \bigcap V \in \tau ,for ~~U,V\in \tau\) iii-\(\Large \cup_{\alpha \in \Delta} V_\alpha \in \tau \) for an indexed family \(\ v_\alpha \in \tau\)

OpenStudy (ikram002p):

also i need you to write down your text book definition for a metric space... the one on wiki is bullsh!t.

OpenStudy (ikram002p):

brb...

OpenStudy (anonymous):

I know some of those, but that does not look familiar at all. Definition: Let (X,d) be a metric space, let \[x_0\in X\] and let \[r>0\] be a positive reel number. Then the set \[\left\{ x \in X|d(x,x_0)<r \right\}\] consisting of the points X which distance to x_0 is less than r, is called the open ball, with center in x_0 and radius r. It is defined: \[B_r(x_0)=\left\{ x \in X|d(x,x_0)<r \right\}\]

Parth (parthkohli):

Waaait, so we really are talking about topological spaces here? Wow that definition is very close to what I've seen.

OpenStudy (anonymous):

Yes, that is correct, topological spaces.

Parth (parthkohli):

D'oh, I could have helped you here then.

OpenStudy (anonymous):

Hmm yea, the definition from my book, looks a lot like the one you found on wikipedia @ParthKohli

OpenStudy (anonymous):

@ganeshie8 Do you know anything about prooving something like this? This is Math-analysis 1, and you've helped med with math-analysis 0 before

ganeshie8 (ganeshie8):

See if this helps http://math.stackexchange.com/questions/1634312/topology-open-set

OpenStudy (anonymous):

That sure does! Thank you :)

ganeshie8 (ganeshie8):

Np :)